Quadtatice model-projectile
I need to find an equation and the initial velocity for the equation-
h(t)=-16t^2+Vo*t+Ho
Ho=is the ball's height (in feet) at moment of release, the initial height
Vo= ball's speed (in feet/sec) at the moment of release, initial velocity
T= time (in seconds) since the ball was released
ho and vo are constants
I threw a ball up about 10 feet
Initial height started: 3.92 feet
Flight time:2.28 seconds
I need help being able to come up with an equation with this data to bring up on my calculator and graph. I don't know how find/solve or how to find the Vo which is the initial velocity.

Answer by stanbon(75887)

Quadtatice model-projectile
I need to find an equation and the initial velocity for the equation-
h(t)=-16t^2+Vo*t+Ho
Ho=is the ball's height (in feet) at moment of release, the initial height
Vo= ball's speed (in feet/sec) at the moment of release, initial velocity
T= time (in seconds) since the ball was released
ho and vo are constants
I threw a ball up about 10 feet
Initial height started: 3.92 feet
Flight time:2.28 seconds
I need help being able to come up with an equation with this data to bring up on my calculator and graph. I don't know how find/solve or how to find the Vo which is the initial velocity.
------
h(0) = 3.92 ft
h(2.28) = 0 ft
h(1.14) = 10 ft
-------------
Equation Form:: -16t^2 + Vot + Ho = h(t)
Use the three equations above to form the following ::
-16*0^2 + Vo*0 + Ho = 3.92
-16*2.28^2 + Vo*2.28 + 3.92= 0
--------------------------------------
Solve the System::
Ho = 3.92
2.28*Vo = 16*2.28^2 - 3.92
2.28*Vo = 79.25
Vo = 34.76 ft/sec
----------------------
Cheers,
Stan H.
--------------